OpenStudy (anonymous):

The length of a field in yards is a function f(n) of the length n in feet. Write a function rule for this situation.

OpenStudy (bahrom7893):

f(n)>0

OpenStudy (bahrom7893):

i don't know what u mean by rule, but since length cannot be negative...

OpenStudy (anonymous):

A.f(n) =1/3n B.f(n) = 3n C.f(n)= 12n D.f(n) = n/12